# A Unified Linear Programming Framework for Offline Reward Learning from Human Demonstrations and Feedback

### v1

@@ 1 Introduction
Reward learning involves inferring and shaping the underlying reward function from observed human demonstrations and feedback. Inverse reinforcement learning (IRL) and reinforcement learning from human feedback (RLHF, also known as preference-based reinforcement learning) are key methodologies in reward learning, applied in various sequential decision-making tasks such as games [1, 2, 3], robotics [4, 5, 6], and language models [7, 8, 9, 10, 11]. Particularly in the recent drastic development of large language models (LLMs), RLHF has played a crucial role in fine-tuning models to better align with human preferences [10]. However, despite the notable empirical success of these algorithms, a significant gap remains in the theoretical analysis of IRL and RLHF, limiting us to guarantee their reliability. This work aims to bridge this gap by proposing a novel theoretical framework for offline IRL and RLHF.
@@
IRL aims to infer a reward function that aligns with an expert behavior from demonstrations [12, 13]. Typical IRL algorithms employ a bi-level optimization framework within the context of maximum likelihood estimation (MLE). In this framework, the inner optimization evaluates the policy based on the current reward parameters, while the outer optimization updates these parameters to better match observed expert behavior. These algorithms have been extensively explored in the literature [14, 15, 16, 17, 18], and their convergence is studied in both online settings [17] and offline settings [18].
@@
Building upon IRL, RLHF recovers a reward function from human feedback data, which is commonly given by pairwise or $K$-wise trajectory comparisons. Despite growing interest in RLHF, only a few algorithms provide optimality guarantees with sample complexity bound. Most existing approaches adopt an MLE framework, assuming that human evaluators follow a presupposed preference model, such as the Bradley-Terry-Luce (BTL) model. The reward parameters are then fine-tuned to maximize the log-likelihood of the collected offline preference data [2]. Recently, [19, 20, 21] have proposed offline RLHF algorithms with optimality guarantees by adopting a pessimistic mechanism from offline reinforcement learning (RL) theory.
@@
While the MLE-based approach offers a solid theoretical framework for both IRL and RLHF, it comes with unavoidable limitations. Particularly in IRL, bi-level optimization algorithms face computational challenges due to their nested-loop structures. In addition, MLE-based algorithms rely on a specific decision-making or preference model they employ. For example, the IRL algorithm proposed by [17] learns the reward function that aligns with an expert policy, which is assumed to be an optimal softmax policy. Furthermore, RLHF algorithms (e.g. [19, 20]) assume a preference model for human evaluator, which might not fully capture the complex and diverse nature of real-world human preferences. Consequently, their optimality guarantees might be compromised if there exists a mismatch between actual human preferences and the model in use.
@@
In response to these challenges inherent in MLE frameworks, recent research has shifted focus from estimating a single reward function (under a presupposed model) to recovering a feasible reward set, a set of rewards where the expert policy is (near) optimal. Notably, [22, 23] estimated the feasible reward set from finite-horizon Bellman equations and provided sample complexity bounds associated with estimation errors. However, their algorithm requires a generative model of state transition probabilities. [24] mitigated this requirement by adopting an efficient exploration policy for sampling trajectories, though it remains in an online setting. More recently, and concurrent with our work, [25] introduced the first offline algorithm with a theoretical guarantee. They introduce a pessimistic mechanism to address the issue of non-uniform data coverage, penalizing state-action pairs with low visitation frequency. Nevertheless, these penalty functions are nonlinear and non-convex, resulting in a non-convex reward set. This could limit flexibility for applications, especially when selecting a specific reward function within the set.
@@
In line with this evolving perspective, we aim to obtain a convex estimate of a feasible reward set in an offline setting. To achieve this, we leverage recent advancements in the Linear Programming (LP) framework within the domain of offline RL. A fundamental challenge in offline RL is the non-uniform coverage of offline data  [26, 27]. To address this issue, recent algorithms have employed a pessimistic mechanism that conservatively selects the value function or model within an uncertainty set [28, 29, 30, 31, 32, 33]. However, these pessimistic approaches often introduce non-convex optimization problems, which can be intractable. In the latest research, a series of works [34, 35, 36] have introduced LP-based methods that relax data coverage assumptions and provide tractable algorithms suitable for function approximation by introducing convex formulations.
@@
  • In Theorem 1, the optimality of the estimated reward set is provided such that any reward function within this set ensures the expert policy is $\tilde{O}(\sqrt{|S||A|/N})$-suboptimal, under appropriate data coverage assumption.
  • In offline RLHF, we align reward functions with pairwise trajectory comparison data using linear constraints (Section 4.1). In Theorem 2, we provide the generalization guarantee of the estimated reward function for unseen trajectory pairs.
@@ 2.1 Markov Decision Process (MDP)
$v^{*}(s)=\max_{a\in A}\left\{r(s,a)+\gamma\sum_{s^{\prime}\in S}P(s^{\prime}|s,a)v^{*}(s^{\prime})\right\},$ (2.1)
which holds for any $s\in S$. It is well-established that $v^{*}$ can be determined by solving the following linear programming (LP) as outlined in [37]:
$\min_{v\in\mathbb{R}^{|S|}}\;(1-\gamma)\mu_{0}^{\top}v\quad\text{s.t.}\quad M^{\top}v-r\geq 0.$ (2.2)
@@
$\max_{d\in\mathbb{R}^{|S||A|}}\;r^{\top}d\quad\text{s.t.}\quad Md=(1-\gamma)\mu_{0},\quad d\geq 0.$ (2.3)
@@
$d^{\pi}(s,a)=(1-\gamma)\sum_{h=0}^{\infty}\gamma^{h}P^{\pi}_{\mu_{0}}(s_{h}=s,a_{h}=a).$ (2.4)
Here, $P^{\pi}_{\mu_{0}}(s_{h}=s,a_{h}=a)$ represents the probability of $(s_{h},a_{h})=(s,a)$, given $s_{0}\sim\mu_{0}$ and $a_{h^{\prime}}\sim\pi(s_{h^{\prime}})$ for all $h^{\prime}\geq 0$. The dependence on $\gamma$ and $\mu_{0}$ is omitted in the notation $d^{\pi}$ for simplicity. A more detailed relationship between $\pi$ and $d^{\pi}$ can be found in [37].
@@ 2.2 Offline IRL: Learning from Expert Trajectories
$\mathcal{D}_{\text{IRL}}=\{\tau^{n}=(s_{0}^{n},a_{0}^{n},s_{1}^{n},\ldots,s_{H-1}^{n},a_{H-1}^{n},s_{H}^{n})\}_{n=1}^{N}.$ (2.5)
@@
$\hat{d}_{e}(s,a)=(1-\gamma)\frac{1}{N}\sum_{h=0}^{H-1}\gamma^{h}N_{h}(s,a)\quad\forall(s,a)\in S\times A.$ (2.6)
@@ 3.1 LP Formulation of Offline IRL
$\begin{split}\max_{d\in\mathbb{R}^{|S||A|}}\;r^{\top}d\quad\text{s.t.}\quad Md=(1-\gamma)\mu_{0},\quad d\geq 0.\end{split}$ (3.1)
@@
In the offline setting, recovering the ground-truth $\mathcal{R}$ is challenging since we only have access to the empirical estimate of $d_{e}$ and the transition matrix $P$. Consequently, our goal is to recover an estimate of $\mathcal{R}$ in which $d_{e}$ is a near-optimal solution to (3.1) for any $r$ in this estimated set.
@@
The primary challenge in offline RL and IRL is the non-uniform coverage of the offline dataset. To address this issue, we adopt the marginal importance sampling (MIS) framework in the literature [38, 39], which considers the scaled version of LP. First, we define the optimization variable $w_{d}\in\mathbb{R}^{|S||A|}$ as
$w_{d}(s,a):=\begin{cases}\frac{d(s,a)}{d_{e}(s,a)}\quad&\text{if}\quad d_{e}(s,a)>0,\\ 0\quad&\text{if}\quad d_{e}(s,a)=0.\end{cases}$ (3.3)
@@
$\begin{split}u(s,a)&:=r(s,a)d_{e}(s,a),\\ K(s^{\prime},(s,a))&:=d_{e}(s,a)\mathbf{1}_{\{s=s^{\prime}\}}-\gamma d^{\prime}_{e}(s,a,s^{\prime}),\end{split}$ (3.4)
where $d^{\prime}_{e}(s,a,s^{\prime}):=d_{e}(s,a)P(s^{\prime}|s,a)$ for any $(s,a,s^{\prime})$. In this MIS framework, $u$ and $K$ correspond to $r$ and $P$, respectively. The following lemma shows this relationship clearly (see Lemma 1 in [36] for the proof).
Lemma 1.
@@
By Lemma 1, the dual LP can be written with $u$, $w$, and $K$ as follows:
$\max_{w\in\mathbb{R}^{|S||A|}}\;u^{\top}w\quad\text{s.t.}\quad Kw=(1-\gamma)\mu_{0},\quad w\geq 0.$ (3.5)
@@
$K_{\mathcal{D}}(s^{\prime},(s,a)):=\hat{d}_{e}(s,a)\mathbf{1}_{\{s=s^{\prime}\}}-\gamma\hat{d}^{\prime}_{e}(s,a,s^{\prime}),$ (3.6)
@@
$\hat{d}^{\prime}_{e}(s,a,s^{\prime}):=(1-\gamma)\frac{1}{N}\sum_{h=0}^{H-1}\gamma^{h}N_{h}(s,a,s^{\prime}).$ (3.7)
@@
$\begin{split}\max_{w\in\mathbb{R}^{|S||A|}}\quad&u_{\mathcal{D}}^{\top}w\\ \text{s.t.}\quad&X^{\top}(K_{\mathcal{D}}w-(1-\gamma)\mu_{0})\leq\epsilon_{x},\quad w\geq 0.\end{split}$ (3.8)
Additionally, the dual of (3.8) can be expressed as
$\begin{split}\min_{v\in\mathbb{R}^{N_{x}}}\quad&(1-\gamma)\mu_{0}^{\top}Xv+\epsilon_{x}^{\top}v\\ \text{s.t.}\quad&K_{\mathcal{D}}^{\top}Xv\geq u_{\mathcal{D}},\quad v\geq 0,\end{split}$ (3.9)
@@
Under the empirical LP formulations, our goal is to estimate the set of $u$ such that $w_{e}$ is (near) optimal to (3.8). Consider the primal-dual optimality conditions of $(v,w)$ under a reward function $u$:
(Primal $\begin{split}&\text{(Primal feasibility)}:K_{\mathcal{D}}^{\top}Xv\geq u,\;v\geq 0,\\ &\text{(Dual feasibility)}:X^{\top}(K_{\mathcal{D}}w-(1-\gamma)\mu_{0})\leq\epsilon_{x},\;w\geq 0,\\ &\text{(Zero duality gap)}:(1-\gamma)\mu_{0}^{\top}Xv+\epsilon_{x}^{\top}v=u^{\top}w.\end{split}$ (3.10)
@@
Consequently, the feasible reward set can be estimated by identifying $(u,v)$ pairs for which $(u,v,w_{e})$ satisfies (3.10). Here, we further relax the zero duality gap condition with the slack parameter $\epsilon_{g}\geq 0$ for two reasons. First, the true reward might not satisfy this equality constraint due to errors in empirical estimation of $d_{e}$ and $K$. Second, we are also interested in the reward such that $\pi_{e}$ is near-optimal. Therefore, we consider the following polyhedron as an estimate of the feasible reward set:
^$\hat{\mathcal{R}}_{\text{IRL}}(\epsilon_{g}):=\{(u,v)\mid\underbrace{(1-\gamma)\mu_{0}^{\top}Xv+\epsilon_{x}^{\top}v-u^{\top}\mathbf{1}\leq\epsilon_{g}}_{(i)},\;\underbrace{K_{\mathcal{D}}^{\top}Xv\geq u,\;v\geq 0}_{(ii)},\;\underbrace{-\hat{d}_{e}\leq u\leq\hat{d}_{e}}_{(iii)}\}.$ (3.11)

### v3

@@ 1 Introduction
Reward learning involves inferring and shaping the underlying reward function from observed human demonstrations and feedback. Inverse reinforcement learning (IRL) and reinforcement learning from human feedback (RLHF, also known as preference-based reinforcement learning) are key methodologies in reward learning, applied in various sequential decision-making tasks such as games (MacGlashan et al., 2017; Christiano et al., 2017; Ibarz et al., 2018), robotics (Finn et al., 2016; Brown et al., 2019; Shin et al., 2023), and language models (Ziegler et al., 2019; Stiennon et al., 2020; Wu et al., 2021; Ouyang et al., 2022; Liu et al., 2023).s Particularly in the recent drastic development of large language models (LLMs), RLHF has played a crucial role in fine-tuning models to better align with human preferences (Ouyang et al., 2022). However, despite the notable empirical success of these algorithms, a significant gap remains in the theoretical analysis of IRL and RLHF, limiting us to guarantee their reliability. This work aims to bridge this gap by proposing a novel theoretical framework for offline IRL and RLHF.
@@
IRL aims to infer a reward function that aligns with an expert behavior from demonstrations (Ng & Russell, 2000; Abbeel & Ng, 2004). Typical IRL algorithms employ a bi-level optimization framework within the context of maximum likelihood estimation (MLE). In this framework, the inner optimization evaluates the policy based on the current reward parameters, while the outer optimization updates these parameters to better match observed expert behavior. These algorithms have been extensively explored in the literature (Ziebart et al., 2008; Wulfmeier et al., 2015; Zhou et al., 2017; Zeng et al., 2022, 2023), and their convergence is studied in both online settings (Zeng et al., 2022) and offline settings (Zeng et al., 2023).
@@
Building upon IRL, RLHF recovers a reward function from human feedback data, which is commonly given by pairwise or $K$-wise trajectory comparisons. Despite growing interest in RLHF, only a few algorithms provide optimality guarantees with sample complexity bound. Most existing approaches adopt an MLE framework, assuming that human evaluators follow a presupposed preference model, such as the Bradley-Terry-Luce (BTL) model. The reward parameters are then fine-tuned to maximize the log-likelihood of the collected offline preference data (Christiano et al., 2017). Recently, (Zhu et al., 2023; Zhan et al., 2023; Li et al., 2023) have proposed offline RLHF algorithms with optimality guarantees by adopting a pessimistic mechanism from offline reinforcement learning (RL) theory. (Park et al., 2024) further suggested a personalized setting that learns multiple reward models for heterogeneous preferences.
@@
While the MLE-based approach offers a solid theoretical framework for both IRL and RLHF, it comes with unavoidable limitations. Particularly in IRL, bi-level optimization algorithms face computational challenges due to their nested-loop structures. In addition, MLE-based algorithms rely on a specific decision-making or preference model they employ. For example, the IRL algorithm proposed by (Zeng et al., 2022) learns the reward function that aligns with an expert policy, which is assumed to be an optimal softmax policy. Furthermore, RLHF algorithms (e.g. (Zhu et al., 2023; Zhan et al., 2023)) assume a preference model for human evaluator, which might not fully capture the complex and diverse nature of real-world human preferences. Consequently, their optimality guarantees might be compromised if there exists a mismatch between actual human preferences and the model in use.
@@
In response to these challenges inherent in MLE frameworks, recent research has shifted focus from estimating a single reward function (under a presupposed model) to recovering a feasible reward set, a set of rewards where the expert policy is (near) optimal. Notably, (Metelli et al., 2021, 2023) estimated the feasible reward set from finite-horizon Bellman equations and provided sample complexity bounds associated with estimation errors. However, their algorithm requires a generative model of state transition probabilities. (Lindner et al., 2022) mitigated this requirement by adopting an efficient exploration policy for sampling trajectories, though it remains in an online setting. More recently, and concurrent with our work, (Zhao et al., 2023) introduced the first offline algorithm with a theoretical guarantee. They introduce a pessimistic mechanism to address the issue of non-uniform data coverage, penalizing state-action pairs with low visitation frequency. Nevertheless, these penalty functions are nonlinear and non-convex, resulting in a non-convex reward set. This could limit flexibility for applications, especially when selecting a specific reward function within the set.
@@
In line with this evolving perspective, we aim to obtain a convex estimate of a feasible reward set in an offline setting. To achieve this, we leverage recent advancements in the Linear Programming (LP) framework within the domain of offline RL. A fundamental challenge in offline RL is the non-uniform coverage of offline data  (Fujimoto et al., 2019; Kumar et al., 2020). To address this issue, recent algorithms have employed a pessimistic mechanism that conservatively selects the value function or model within an uncertainty set (Liu et al., 2020; Jin et al., 2021; Rashidinejad et al., 2021; Xie et al., 2021; Uehara & Sun, 2021; Chen & Jiang, 2022). However, these pessimistic approaches often introduce non-convex optimization problems, which can be intractable. In the latest research, a series of works (Zhan et al., 2022; Rashidinejad et al., 2022; Ozdaglar et al., 2023) have introduced LP-based methods that relax data coverage assumptions and provide tractable algorithms suitable for function approximation by introducing convex formulations.
@@
  • In Theorem 3.3, the optimality of the estimated reward set is provided such that any reward function within this set ensures the expert policy is $\tilde{O}(\sqrt{|S||A|/N})$-suboptimal, under appropriate data coverage assumption.
  • In offline RLHF, we align reward functions with pairwise trajectory comparison data using linear constraints (Section 4.1). In Theorem 4.1, we provide the generalization guarantee of the estimated reward function for unseen trajectory pairs.
@@ 2.1 Markov Decision Process (MDP)
$v^{*}(s)=\max_{a\in A}\left\{r(s,a)+\gamma\sum_{s^{\prime}\in S}P(s^{\prime}|s,a)v^{*}(s^{\prime})\right\},$ (1)
which holds for any $s\in S$. It is well-established that $v^{*}$ can be determined by solving the following linear programming (LP) as outlined in (Puterman, 2014):
$\min_{v\in\mathbb{R}^{|S|}}\;(1-\gamma)\mu_{0}^{\top}v\quad\text{s.t.}\quad M^{\top}v-r\geq 0.$ (2)
@@
$\max_{d\in\mathbb{R}^{|S||A|}}\;r^{\top}d\quad\text{s.t.}\quad Md=(1-\gamma)\mu_{0},\quad d\geq 0.$ (3)
@@
$d^{\pi}(s,a)=(1-\gamma)\sum_{h=0}^{\infty}\gamma^{h}P^{\pi}_{\mu_{0}}(s_{h}=s,a_{h}=a).$ (4)
Here, $P^{\pi}_{\mu_{0}}(s_{h}=s,a_{h}=a)$ represents the probability of $(s_{h},a_{h})=(s,a)$, given $s_{0}\sim\mu_{0}$ and $a_{h^{\prime}}\sim\pi(s_{h^{\prime}})$ for all $h^{\prime}\geq 0$. The dependence on $\gamma$ and $\mu_{0}$ is omitted in the notation $d^{\pi}$ for simplicity. A more detailed relationship between $\pi$ and $d^{\pi}$ can be found in (Puterman, 2014).
@@ 2.2 Offline IRL: Learning from Expert Trajectories
$\mathcal{D}_{\text{IRL}}=\{\tau^{n}=(s_{0}^{n},a_{0}^{n},s_{1}^{n},\ldots,s_{H-1}^{n},a_{H-1}^{n},s_{H}^{n})\}_{n=1}^{N}.$ (5)
@@
$\hat{d}_{e}(s,a)=(1-\gamma)\frac{1}{N}\sum_{h=0}^{H-1}\gamma^{h}N_{h}(s,a)\quad\forall(s,a)\in S\times A.$ (6)
@@ 3.1 LP Formulation of Offline IRL
$\begin{split}\max_{d\in\mathbb{R}^{|S||A|}}\;r^{\top}d\quad\text{s.t.}\quad Md=(1-\gamma)\mu_{0},\quad d\geq 0.\end{split}$ (7)
@@
In the offline setting, recovering the ground-truth $\mathcal{R}$ is challenging since we only have access to the empirical estimate of $d_{e}$ and the transition matrix $P$. Consequently, our goal is to recover an estimate of $\mathcal{R}$ in which $d_{e}$ is a near-optimal solution to (7) for any $r$ in this estimated set.
@@
The primary challenge in offline RL and IRL is the non-uniform coverage of the offline dataset. To address this issue, we adopt the marginal importance sampling (MIS) framework in the literature (Nachum et al., 2019; Lee et al., 2021), which considers the scaled version of LP. First, we define the optimization variable $w_{d}\in\mathbb{R}^{|S||A|}$ as
$w_{d}(s,a):=\begin{cases}\frac{d(s,a)}{d_{e}(s,a)}\quad&\text{if}\quad d_{e}(s,a)>0,\\ 0\quad&\text{if}\quad d_{e}(s,a)=0.\end{cases}$ (9)
@@
$\begin{split}u(s,a)&:=r(s,a)d_{e}(s,a),\\ K(s^{\prime},(s,a))&:=d_{e}(s,a)\mathbf{1}_{\{s=s^{\prime}\}}-\gamma d^{\prime}_{e}(s,a,s^{\prime}),\end{split}$ (10)
where $d^{\prime}_{e}(s,a,s^{\prime}):=d_{e}(s,a)P(s^{\prime}|s,a)$ for any $(s,a,s^{\prime})$. In this MIS framework, $u$ and $K$ correspond to $r$ and $P$, respectively. The following lemma shows this relationship clearly (see Lemma 1 in (Ozdaglar et al., 2023) for the proof).
Lemma 3.1.
@@
By Lemma 3.1, the dual LP can be written with $u$, $w$, and $K$ as follows:
$\max_{w\in\mathbb{R}^{|S||A|}}\;u^{\top}w\quad\text{s.t.}\quad Kw=(1-\gamma)\mu_{0},\quad w\geq 0.$ (11)
@@
$K_{\mathcal{D}}(s^{\prime},(s,a)):=\hat{d}_{e}(s,a)\mathbf{1}_{\{s=s^{\prime}\}}-\gamma\hat{d}^{\prime}_{e}(s,a,s^{\prime}),$ (12)
@@
$\hat{d}^{\prime}_{e}(s,a,s^{\prime}):=(1-\gamma)\frac{1}{N}\sum_{h=0}^{H-1}\gamma^{h}N_{h}(s,a,s^{\prime}).$ (13)
@@
$\begin{split}\max_{w\in\mathbb{R}^{|S||A|}}\quad&u_{\mathcal{D}}^{\top}w\\ \text{s.t.}\quad&X^{\top}(K_{\mathcal{D}}w-(1-\gamma)\mu_{0})\leq\epsilon_{x},\quad w\geq 0.\end{split}$ (14)
Additionally, the dual of (14) can be expressed as
$\begin{split}\min_{v\in\mathbb{R}^{N_{x}}}\quad&(1-\gamma)\mu_{0}^{\top}Xv+\epsilon_{x}^{\top}v\\ \text{s.t.}\quad&K_{\mathcal{D}}^{\top}Xv\geq u_{\mathcal{D}},\quad v\geq 0,\end{split}$ (15)
@@
Under the empirical LP formulations, our goal is to estimate the set of $u$ such that $w_{e}$ is (near) optimal to (14). Consider the primal-dual optimality conditions of $(v,w)$ under a reward function $u$:
(Primal $\begin{split}&\text{(Primal feasibility)}:K_{\mathcal{D}}^{\top}Xv\geq u,\;v\geq 0,\\ &\text{(Dual feasibility)}:X^{\top}(K_{\mathcal{D}}w-(1-\gamma)\mu_{0})\leq\epsilon_{x},\;w\geq 0,\\ &\text{(Zero duality gap)}:(1-\gamma)\mu_{0}^{\top}Xv+\epsilon_{x}^{\top}v=u^{\top}w.\end{split}$ (16)
@@
Consequently, the feasible reward set can be estimated by identifying $(u,v)$ pairs for which $(u,v,w_{e})$ satisfies (16). Here, we further relax the zero duality gap condition with the slack parameter $\epsilon_{g}\geq 0$ for two reasons. First, the true reward might not satisfy this equality constraint due to errors in empirical estimation of $d_{e}$ and $K$. Second, we are also interested in the reward such that $\pi_{e}$ is near-optimal. Therefore, we consider the following polyhedron as an estimate of the feasible reward set:
^$\begin{split}\hat{\mathcal{R}}_{\text{IRL}}(\epsilon_{g}):=\{(u,v)\mid\underbrace{(1-\gamma)\mu_{0}^{\top}Xv+\epsilon_{x}^{\top}v-u^{\top}\mathbf{1}\leq\epsilon_{g}}_{(i)},\\ \underbrace{K_{\mathcal{D}}^{\top}Xv\geq u,\;v\geq 0}_{(ii)},\;\underbrace{-\hat{d}_{e}\leq u\leq\hat{d}_{e}}_{(iii)}\}.\end{split}$ (17)